OpenStudy (anonymous):

An 80.0-gram sample of a gas was heated from 25 °C to 225 °C. During this process, 346 J of work was done by the system and its internal energy increased by 7845 J. What is the specific heat of the gas?

OpenStudy (aaronq):

q=mCdT dT = T2-T1= 225-25 Celcius m=80 g q=7845 J C = specific heat capacity (what you're solving for)